OpenStudy (osameh93):

Find the inverse of F(X)= 3x²+5x-2 , x>=0

OpenStudy (anonymous):

F(x)/3 = x^2 + 5x/3 - 2/3 (F(x)+2)/3 = x^2 + 5x/3 5/3 = 2ab a=1 b = 5/6 b^2 = 25/36 (F(x)+2)/3 = (x + 5/6)^2 - 25/36 (F(x)+2)/3 + 25/36 = (x + 5/6)^2 sqrt((F(x)+2)/3 + 25/36) -5/6 = x sqrt((x+2)/3 + 25/36) -5/6 = y